# Primordial gravitational waves in Wheeler-DeWitt non-commutative linearized branch-cut quantum gravity

### 12 Jan 2024 v1

@@ VI Stochastic gravitational wave background
VI.1 Sub-horizon condition with a matter-energy source
@@
Combining equations (86), (93), and (94), the sub-horizon condition leads to the differential equation
${h^{\prime\prime}}(k,\tau)+k^{2}{h}(k,\tau)=-\frac{16}{3}\pi\eta\Biggl{(}-\frac{3\alpha\gamma}{\bigl{(}3\alpha-1\bigr{)}\eta^{3\alpha-1}}+\Biggl{[}g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}\Biggr{]}\tau-\frac{3}{2}\frac{\alpha^{2}\tau^{2}}{\eta^{6\alpha-1}}\Biggr{)}\,.$ (115)
The aforementioned field equation describes the evolution over time of the metric disturbance having as its source the potential defined in equation (93) that configures the dynamic composition of matter in the primordial cosmic period. To solve equation (115), due to the technical difficulties of its resolution, we used powerful computational algorithms, based on the Runge-Kutta-Fehlberg method, that made it possible to obtain algebraic solutions without the need to implement simplifications or any computational approximation (see Appendix C).
VI.2 Generation of gravitational waves: bubble collisions
@@
$h_{0}^{2}\Omega_{b}(\nu/\nu_{b})=1.67\times 10^{-5}\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{2}\Biggl{(}\frac{\kappa_{\infty}\alpha_{T}}{1+\alpha_{T}}\Biggr{)}^{2}\Bigl{(}\frac{100}{g_{*}}\Bigr{)}^{1/3}\Biggl{(}\frac{0.11u^{3}_{\omega}}{0.42+u^{2}_{\omega}}\Biggr{)}S_{b}(\nu/\nu_{b})\,,$ (116)
@@
$\nu_{b}=1.65$$\times$ $10^{-5}\Bigl{(}\frac{\beta}{H_{*}}\Bigr{)}\Bigl{(}\frac{T_{*}}{100~{}{\rm GeV}}\Bigr{)}\Bigl{(}\frac{g_{*}}{100}\Bigr{)}$
$\times$ $\Biggl{(}\frac{0.62}{1.8+u^{2}_{\omega}-0.1u_{\omega}}\Biggr{)}Hz\approx 1{\rm mHz}\,.$
@@
In the following, due to the formal complexity of the solutions of equation (115), guided by the dominant terms of these solutions in the domain $\nu/\nu_{c}\to 0$ (see Appendix C), we propose in expressions (118) and (119) an analytical formulation that maintains the essence of the branch-cut quantum gravity proposal as well as the main requirements of the spectral shape of the signal. Combining equation (114) and the solution of (115), by means of a systematic preliminary analytical study summarized in Appendix C, we propose the following expression to represent the spectral shape of the signal:
$S_{b}(\nu/\nu_{c})=\sum_{m=0}^{\kappa}(-1)^{m+1}\frac{b}{a^{m}\bigl{(}\frac{\nu}{\nu_{c}}\bigr{)}^{mb-a}}\,,$ (118)
@@
$S_{b}(\nu/\nu_{c})\Rightarrow\frac{b\times\bigl{(}\frac{\nu}{\nu_{b}}\bigr{)}^{a}}{a\times\bigl{(}\frac{\nu}{\nu_{b}}\bigr{)}^{b}+1}+{\cal O}\Bigl{(}\Bigl{(}\frac{\nu}{\nu_{b}}\Bigr{)}^{\kappa b-a}\Bigr{)}\,,$ (119)
in order $m=\kappa$. As for the content of the initial proposal for the branched quantum gravity, by introducing a potential that would describe the main content of the relic Universe in terms of mass and energy contributions, against the backdrop of a non-commutative algebraic structure, sources of the primordial gravitational waves, these aspects are also well covered in the proposal considered. After carrying out the different theoretical steps, we arrive to a formal dependence on the corresponding spectral shape of the signal in terms of a convergent series $(\nu/\nu_{c})^{\kappa}$-dependent, in different orders. In short, our theoretical developments indicate that the final expression for the spectral shape of the signal should reduce to a regular function with may be expanded in $1/(\nu/\nu_{c})^{\kappa}$-dependent series. In addition, the proposal defined in equation (119) contemplates the condition that the maximum of the function corresponds to the value of the characteristic frequency, $\nu/\nu_{c}=1$, as well as additionally makes it possible to fit simulated data with the simple choice of the only formulation parameter, $a$ (see for example [5], in which the choice of $a$ corresponds to $2.8$ to fit simulated data to describe spectrum of the gravitational radiation determined by numerical simulations using the ’envelope approximation’ [78]).
VII Discussion of Results
There are numerous research works in the literature that address the topic of quantum gravity in a non-commutative environment. However, with regard to the description of the wave function of the Universe based on the formalism of Hořava-Lifshitz quantum gravity and the Wheeler-DeWitt equation, or based on other alternative approaches, the number of articles dealing with this topic are not significant. Most of the published works deal with the temporal evolution of the scale factor of the Universe, a topic of studies commonly designated as “dynamical equations”. In some articles the authors deal with both topics. We indicate a few references on these topics [79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90]). We emphasize once again that concerning the wave function of the Universe in a non-commutative environment, most authors, due to the inherent computational difficulties imposed by the formalism, use approximations that significantly limit their conclusions and direct their studies towards formal aspects without a numerical approach or algebraic results.
@@ VIII Conclusions
$\Delta S\Delta E\lesssim\frac{\hbar c}{2}\frac{\pi k_{B}\bigl{(}1+\theta\bigr{)}}{\sqrt{\theta}\ell_{P}}\,,$ (122)
@@
In this contribution, the results obtained for the description of gravitational waves bring a new perspective, in line with the implications of equation (122) which indicate an uncertainty around the previous boundary conditions, as a result of a minimum scale. And also in line with the implications of the Bekenstein criterion, indicating a quantum leap in the boundary of the branch-cut transition region. These results also indicate, in view of the mixture of intensities of the original sources of relic gravitational waves identified by the distribution of colors associated with the quantities that describe the corresponding power spectra and density parameters, two fundamental aspects. The first concerns the primordial vision of a smooth and isotropic Universe, symmetrical to ours. In this description, the stress-energy must take the form of a perfect fluid fully described by an energy density $\rho_{\epsilon}$ and pressure p, in which the variations of the internal equation of state are so small that they make no qualitative difference in the cosmic evolutionary process. This premise seems to collapse when we examine the results obtained in this contribution, which indicate asymmetric Universes, whether in the distribution of intensities of gravitational wave sources, or in the implications of non-symmetric algebra in the short-distance structure of spacetime. The second aspect concerns the transition region between the different phases of the Universe, bringing to light the discussion about the possibility of observing signatures not only corresponding to the era marked by the dissociation of matter and radiation in our Universe, but also the phases associated with the cosmic seeds, sources of gravitational waves, existing in the mirror Universe.
@@
A final word regarding the results obtained to describe relic gravitational waves in a non-commutative formulation: the regular function that describes the spectral shape of the signal is the result, among many other functional possibilities, of the Lagrangian structure of the branched quantum gravity formulation. proposal. There are still aspects to be revealed in the future, in order to introduce a functional formulation, albeit regular, that highlights the behavior of the spectral form of the signal parametrically dependent on the non-commutative structure. This aspect, although in our formulation it is implicitly contained in the functional structure of the expression (119), deserves a consistent specific approach.
IX Acknowledgements
@@
$({\tilde{x}}_{i})=({\tilde{u}},\tilde{p}_{u},{\tilde{v}},{\tilde{p}}_{v})\quad\mbox{and}\quad(x_{i})=(u,p_{u},v,p_{v})~{}.$ (123)
@@
$\left\{{\tilde{x}}_{i},{\tilde{x}}_{j}\right\}={\tilde{g}}_{ij}\quad\mbox{and}\quad\left\{x_{i},x_{j}\right\}=g_{ij}\,,$ (124)
@@
${\tilde{g}}_{ji}=-{\tilde{g}}_{ij}\quad\mbox{and}\quad g_{ji}=-g_{ij}~{}.$ (125)
@@
$\left({\tilde{g}}\right)$ $=$ $\left(\begin{array}[]{cccc}0&1&0&0\\ -1&0&0&0\\ 0&0&0&1\\ 0&0&-1&0\end{array}\right)\,;$ (130)
$\left(g\right)$ $=$ $\left(\begin{array}[]{cccc}0&1&0&\gamma\\ -1&0&-\chi&\alpha\\ 0&\chi&0&1\\ -\gamma&-\alpha&-1&0\end{array}\right)~{}.$ (135)
Using (124), this leads to the non-zero Poisson brackets (only the Poisson brackets are listed which will be non-zero in the non-commuting case)
@@
$\left\{{\tilde{v}},{\tilde{p}}_{u}\right\}$ $=$ $0\,;\quad\mbox{and}\quad\left\{{\tilde{p}}_{u},{\tilde{p}}_{v}\right\}=0\,;$ (136)
@@
$\left\{v,p_{u}\right\}$ $=$ $\chi\,;\quad\mbox{and}\quad\left\{p_{u},p_{v}\right\}=\alpha~{}.$ (137)
@@
${\tilde{x}}_{i}$ $=$ $\sum_{j}M_{ij}x_{j}\,,$ (138)
such that the above Poisson brackets are satisfied. This gives us conditions for the matrix elements $M_{ij}$. Equation (138) can be cast into the form
@@
$\left(M\right)$ $=$ $\left(\begin{array}[]{cccc}1&0&0&0\\ M_{21}&M_{22}&M_{23}&M_{24}\\ 0&0&1&0\\ M_{41}&M_{42}&M_{43}&M_{44}\end{array}\right)\,.$ (144)
Using (137), and the transformation (138) with (144), we obtain
@@
$\left\{{\tilde{v}},{\tilde{p}}_{u}\right\}$ $=$ $0=M_{22}\chi+M_{24}~{},$ (145)
where the Poisson-bracket $\left\{\tilde{p}_{u},\tilde{p}_{v}\right\}=0$ will be calculated later. First of all we resolve the set of equations (145), which leads to
@@
$M_{42}$ $=$ $-\frac{\gamma}{1-\gamma\chi}\,;\quad M_{44}=\frac{1}{1-\gamma\chi}~{},$ (146)
@@
$\left(M\right)$ $=$ $\left(\begin{array}[]{cccc}1&0&0&0\\ M_{21}&\frac{1}{1-\gamma\chi}&M_{23}&-\frac{\chi}{1-\gamma\chi}\\ 0&0&1&0\\ M_{41}&-\frac{\gamma}{1-\gamma\chi}&M_{43}&\frac{1}{1-\gamma\chi}\end{array}\right)~{}.$ (151)
@@
$\left\{{\tilde{p}}_{u},{\tilde{p}}_{v}\right\}=0=-M_{41}+M_{23}+\frac{\alpha}{a-\gamma\chi}~{},$ (152)
@@
$M_{41}$ $=$ $\frac{\alpha}{1-\gamma\chi}+M_{23}~{}~{}~{}.$ (153)
@@
${\tilde{p}}_{u}$ $=$ $M_{21}u+\frac{1}{1-\gamma\chi}p_{u}+M_{23}v-\frac{\chi}{1-\gamma\chi}p_{v}$
${\tilde{p}}_{v}$ $=$ $\left(\frac{\alpha}{1-\gamma\chi}+M_{23}\right)u-\frac{\gamma}{1-\gamma\chi}p_{u}$ (154)
$+$ $M_{43}v+\frac{1}{1-\gamma\chi}p_{v}~{}~{}~{}.$
We also define $\Gamma=(\gamma\chi-1)$.
@@
Appendix B

### 03 Feb 2024 v2

@@ 6 Stochastic gravitational wave background
6.1 Sub-horizon condition with a matter-energy source
@@
Combining equations (5.13), (5.20), and (5.21), the sub-horizon condition leads to the differential equation
${h^{\prime\prime}}(k,\tau)+k^{2}{h}(k,\tau)$ $=$ $-\frac{16}{3}\pi\eta\Biggl{(}-\frac{3\alpha\gamma}{\bigl{(}3\alpha-1\bigr{)}\eta^{3\alpha-1}}-\frac{3}{2}\frac{\alpha^{2}\tau^{2}}{\eta^{6\alpha-1}}$ (6.1)
$+\Biggl{[}g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}\Biggr{]}\tau\Biggr{)}\,.$
The aforementioned field equation describes the evolution over time of the metric disturbance having as its source the potential defined in equation (5.20) that configures the dynamic composition of matter in the primordial cosmic period. To solve equation (6.1), due to the technical difficulties of its resolution, we used powerful computational algorithms, based on the Runge-Kutta-Fehlberg method, that made it possible to obtain algebraic solutions without the need to implement simplifications or any computational approximation (see Appendix C).
6.2 Generation of gravitational waves: bubble collisions
@@
$h_{0}^{2}\Omega_{b}(\nu/\nu_{b})=1.67\times 10^{-5}\Bigl{(}\frac{H_{*}}{\beta}\Bigr{)}^{2}\Biggl{(}\frac{\kappa_{\infty}\alpha_{T}}{1+\alpha_{T}}\Biggr{)}^{2}\Bigl{(}\frac{100}{g_{*}}\Bigr{)}^{1/3}\Biggl{(}\frac{0.11u^{3}_{\omega}}{0.42+u^{2}_{\omega}}\Biggr{)}S_{b}(\nu/\nu_{b})\,,$ (6.2)
@@
$\nu_{b}=1.65\times 10^{-5}\Bigl{(}\frac{\beta}{H_{*}}\Bigr{)}\Bigl{(}\frac{T_{*}}{100~{}{\rm GeV}}\Bigr{)}\Bigl{(}\frac{g_{*}}{100}\Bigr{)}\Biggl{(}\frac{0.62}{1.8+u^{2}_{\omega}-0.1u_{\omega}}\Biggr{)}Hz\approx 1{\rm mHz}\,.$ (6.3)
@@
In the following, due to the formal complexity of the solutions of equation (6.1), guided by the dominant terms of these solutions in the domain $\nu/\nu_{c}\to 0$ (see Appendix C), we propose in expressions (6.4) and (6.5) an analytical formulation that maintains the essence of the branch-cut quantum gravity proposal as well as the main requirements of the spectral shape of the signal. Combining equation (5.41) and the solution of (6.1), by means of a systematic preliminary analytical study summarized in Appendix C, we propose the following expression to represent the spectral shape of the signal:
$S_{b}(\nu/\nu_{c})=\sum_{m=0}^{\kappa}(-1)^{m+1}\frac{b}{a^{m}\bigl{(}\frac{\nu}{\nu_{c}}\bigr{)}^{mb-a}}\,,$ (6.4)
@@
$S_{b}(\nu/\nu_{c})\Rightarrow\frac{b\times\bigl{(}\frac{\nu}{\nu_{b}}\bigr{)}^{a}}{a\times\bigl{(}\frac{\nu}{\nu_{b}}\bigr{)}^{b}+1}+{\cal O}\Bigl{(}\Bigl{(}\frac{\nu}{\nu_{b}}\Bigr{)}^{\kappa b-a}\Bigr{)}\,,$ (6.5)
in order $m=\kappa$. As for the content of the initial proposal for the branched quantum gravity, by introducing a potential that would describe the main content of the relic Universe in terms of mass and energy contributions, against the backdrop of a non-commutative algebraic structure, sources of the primordial gravitational waves, these aspects are also well covered in the proposal considered. After carrying out the different theoretical steps, we arrive to a formal dependence on the corresponding spectral shape of the signal in terms of a convergent series $(\nu/\nu_{c})^{\kappa}$-dependent, in different orders. In short, our theoretical developments indicate that the final expression for the spectral shape of the signal should reduce to a regular function with may be expanded in $1/(\nu/\nu_{c})^{\kappa}$-dependent series. In addition, the proposal defined in equation (6.5) contemplates the condition that the maximum of the function corresponds to the value of the characteristic frequency, $\nu/\nu_{c}=1$, as well as additionally makes it possible to fit simulated data with the simple choice of the only formulation parameter, $a$ (see for example [5], in which the choice of $a$ corresponds to $2.8$ to fit simulated data to describe spectrum of the gravitational radiation determined by numerical simulations using the ’envelope approximation’ [77]).
7 Discussion of Results
There are numerous research works in the literature that address the topic of quantum gravity in a non-commutative environment. However, with regard to the description of the wave function of the Universe based on the formalism of Hořava-Lifshitz quantum gravity and the Wheeler-DeWitt equation, or based on other alternative approaches, the number of articles dealing with this topic are not significant. Most of the published works deal with the temporal evolution of the scale factor of the Universe, a topic of studies commonly designated as “dynamical equations". In some articles the authors deal with both topics. We indicate a few references on these topics [78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89]). We emphasize once again that concerning the wave function of the Universe in a non-commutative environment, most authors, due to the inherent computational difficulties imposed by the formalism, use approximations that significantly limit their conclusions and direct their studies towards formal aspects without a numerical approach or algebraic results.
@@ 8 Conclusions
$\Delta S\Delta E\lesssim\frac{\hbar c}{2}\frac{\pi k_{B}\bigl{(}1+\theta\bigr{)}}{\sqrt{\theta}\ell_{P}}\,,$ (8.3)
@@
In this contribution, the results obtained for the description of gravitational waves bring a new perspective, in line with the implications of equation (8.3) which indicate an uncertainty around the previous boundary conditions, as a result of a minimum scale. And also in line with the implications of the Bekenstein criterion, indicating a quantum leap in the boundary of the branch-cut transition region. These results also indicate, in view of the mixture of intensities of the original sources of relic gravitational waves identified by the distribution of colors associated with the quantities that describe the corresponding power spectra and density parameters, two fundamental aspects. The first concerns the primordial vision of a smooth and isotropic Universe, symmetrical to ours. In this description, the stress-energy must take the form of a perfect fluid fully described by an energy density $\rho_{\epsilon}$ and pressure p, in which the variations of the internal equation of state are so small that they make no qualitative difference in the cosmic evolutionary process. This premise seems to collapse when we examine the results obtained in this contribution, which indicate asymmetric Universes, whether in the distribution of intensities of gravitational wave sources, or in the implications of non-symmetric algebra in the short-distance structure of spacetime. The second aspect concerns the transition region between the different phases of the Universe, bringing to light the discussion about the possibility of observing signatures not only corresponding to the era marked by the dissociation of matter and radiation in our Universe, but also the phases associated with the cosmic seeds, sources of gravitational waves, existing in the mirror Universe.
@@
A final word regarding the results obtained to describe relic gravitational waves in a non-commutative formulation: the regular function that describes the spectral shape of the signal is the result, among many other functional possibilities, of the Lagrangian structure of the branched quantum gravity formulation. proposal. There are still aspects to be revealed in the future, in order to introduce a functional formulation, albeit regular, that highlights the behavior of the spectral form of the signal parametrically dependent on the non-commutative structure. This aspect, although in our formulation it is implicitly contained in the functional structure of the expression (6.5), deserves a consistent specific approach.
9 Acknowledgements
@@ Appendix A Commuting and non-commuting coordinates
$({\tilde{x}}_{i})=({\tilde{u}},\tilde{p}_{u},{\tilde{v}},{\tilde{p}}_{v})\quad\mbox{and}\quad(x_{i})=(u,p_{u},v,p_{v})~{}.$ (A.1)
@@
$\left\{{\tilde{x}}_{i},{\tilde{x}}_{j}\right\}={\tilde{g}}_{ij}\quad\mbox{and}\quad\left\{x_{i},x_{j}\right\}=g_{ij}\,,$ (A.2)
@@
${\tilde{g}}_{ji}=-{\tilde{g}}_{ij}\quad\mbox{and}\quad g_{ji}=-g_{ij}~{}.$ (A.3)
@@
$\left({\tilde{g}}\right)=\left(\begin{array}[]{cccc}0&1&0&0\\ -1&0&0&0\\ 0&0&0&1\\ 0&0&-1&0\end{array}\right)\,;\quad\left(g\right)=\left(\begin{array}[]{cccc}0&1&0&\gamma\\ -1&0&-\chi&\alpha\\ 0&\chi&0&1\\ -\gamma&-\alpha&-1&0\end{array}\right)~{}.$ (A.12)
Using (A.2), this leads to the non-zero Poisson brackets (only the Poisson brackets are listed which will be non-zero in the non-commuting case)
@@
$\left\{{\tilde{v}},{\tilde{p}}_{u}\right\}$ $=$ $0\,;\quad\mbox{and}\quad\left\{{\tilde{p}}_{u},{\tilde{p}}_{v}\right\}=0\,;$ (A.13)
@@
$\left\{v,p_{u}\right\}$ $=$ $\chi\,;\quad\mbox{and}\quad\left\{p_{u},p_{v}\right\}=\alpha~{}.$ (A.14)
@@
${\tilde{x}}_{i}$ $=$ $\sum_{j}M_{ij}x_{j}\,,$ (A.15)
such that the above Poisson brackets are satisfied. This gives us conditions for the matrix elements $M_{ij}$. Equation (A.15) can be cast into the form
@@
$\left(M\right)$ $=$ $\left(\begin{array}[]{cccc}1&0&0&0\\ M_{21}&M_{22}&M_{23}&M_{24}\\ 0&0&1&0\\ M_{41}&M_{42}&M_{43}&M_{44}\end{array}\right)\,.$ (A.21)
Using (A.14), and the transformation (A.15) with (A.21), we obtain
@@
$\left\{{\tilde{v}},{\tilde{p}}_{u}\right\}$ $=$ $0=M_{22}\chi+M_{24}~{},$ (A.22)
where the Poisson-bracket $\left\{\tilde{p}_{u},\tilde{p}_{v}\right\}=0$ will be calculated later. First of all we resolve the set of equations (A.22), which leads to
@@
$M_{42}$ $=$ $-\frac{\gamma}{1-\gamma\chi}\,;\quad M_{44}=\frac{1}{1-\gamma\chi}~{},$ (A.23)
@@
$\left(M\right)$ $=$ $\left(\begin{array}[]{cccc}1&0&0&0\\ M_{21}&\frac{1}{1-\gamma\chi}&M_{23}&-\frac{\chi}{1-\gamma\chi}\\ 0&0&1&0\\ M_{41}&-\frac{\gamma}{1-\gamma\chi}&M_{43}&\frac{1}{1-\gamma\chi}\end{array}\right)~{}.$ (A.28)
@@
$\left\{{\tilde{p}}_{u},{\tilde{p}}_{v}\right\}=0=-M_{41}+M_{23}+\frac{\alpha}{a-\gamma\chi}~{},$ (A.29)
@@
$M_{41}$ $=$ $\frac{\alpha}{1-\gamma\chi}+M_{23}~{}~{}~{}.$ (A.30)
@@
${\tilde{p}}_{u}$ $=$ $M_{21}u+\frac{1}{1-\gamma\chi}p_{u}+M_{23}v-\frac{\chi}{1-\gamma\chi}p_{v}~{};$
${\tilde{p}}_{v}$ $=$ $\left(\frac{\alpha}{1-\gamma\chi}+M_{23}\right)u-\frac{\gamma}{1-\gamma\chi}p_{u}+M_{43}v+\frac{1}{1-\gamma\chi}p_{v}~{},$ (A.31)
where we have defined $\Gamma\equiv(\gamma\chi-1)$.
@@
Appendix B Generation of gravitational waves: bubble collisions